A genetic experiment involving peas yielded one sample of offspring consisting of 435 green peas and 133 yellow peas. Use a 0.01 significance level to test the claim that under the same​ circumstances, 27 ​% of offspring peas will be yellow. Identify the null​ hypothesis, alternative​ hypothesis, test​ statistic, P-value, conclusion about the null​ hypothesis, and final conclusion that addresses the original claim. Use the​ P-value method and the normal distribution as an approximation to the binomial distribution. What are the null and alternative​ hypotheses? A. Upper H 0 : p not equals 0.27 Upper H 1 : p less than 0.27 B. Upper H 0 : p not equals 0.27 Upper H 1 : p greater than 0.27 C. Upper H 0 : p equals 0.27 Upper H 1 : p greater than 0.27 D. Upper H 0 : p equals 0.27 Upper H 1 : p not equals 0.27 E. Upper H 0 : p not equals 0.27 Upper H 1 : p equals 0.27 F. Upper H 0 : p equals 0.27 Upper H 1 : p less than 0.27 What is the test​ statistic? zequals nothing ​(Round to two decimal places as​ needed.) What is the​ P-value? ​P-valueequals nothing ​(Round to four decimal places as​ needed.) What is the conclusion about the null​ hypothesis? A. Fail to reject the null hypothesis because the​ P-value is greater than the significance​ level, alpha
Answer #1

Answer)

Ho : P = 0.27

Ha : P is not equal to 0.27

N = 435

P = 0.27

First we need to check the conditions of normality that is if n*p and n*(1-p) both are greater than 5 or not

N*p = 117.45

N*(1-p) = 317.55

Both the conditions are met so we can use standard normal z table to estimate the P-Value

Test statistics z = (oberved p - claimed p)/standard error

Standard error = √{claimed p*(1-claimed p)/√n

Observed p = 133/435

Claimed p = 0.27

N = 435

Z = 1.68

From z table,p(z>1.68) = 0.0465

But our test is two tailed

So, P-Value is = 2*0.0465 = 0.093

As the obtained P-Value is > 0.01 (given significance level)

We fail to reject Ho

Fail to reject the null hypothesis because the​ P-value is greater than the significance​ level, alpha